Explain fully the concept of “Twin Deficits” and use this concept to explain why the cause of the US trade deficit with China is largely due to American wars across the globe.
Answer #1

The " Twin Deficits " hypothesis tells us that there is a strong linkage between a nation's government budget balance and current account balance .

Let us see the proof of this theory :

When ( Tax - Government expenditure ) or ( T - G ) is negative we have a budget deficit .

We know , Y = C + I + G + NX

S ( savings ) = Y - C - T

So, S = I + G + NX - T

Therefore , NX = ( S - I ) + ( T - G ) [ ( S - I ) = private saving - investment ]

NX ( current account ) = ( S - I ) + budget balance

This equation proves that current account will fall when there is a budget deficit . ( G > T )

American war across the globe causes government expenditure to increase than revenue , this is because a war means more supplies required and the government has to maintain the army . Thus , this compels US to import more goods and services from foreign countries . Also production and exports fall due to resources used up in war . This results in trade deficit with China . US imports more from China and exports a fewer goods . NX ( Export - Import )